Tetraethyllead, the "lead" additive in leaded gasoline, has a percent composition by mass of 64.06% Pb, 29.71% C and 6.23% H. What is the empirical formula of tetraethyllead?
Answer #1

the empirical formula of a compound can be calculated using the following method

element mass % atomic weight relative number of atom simplest ratio simplest whole number ratio
C 29.71 12 29.71/12=2.48 2.48/0.31=8 8
H 6.23 1 6.23/1=6.23 6.23/0.31=20 20
Pb 64.06 207 64.06/207=0.31 0.31/0.31=1 1

ratio between Pb:C:H =1:8:20

emperical formula containg Pb.C and H is PbC8H20